# Have We Scene It All? Scene Graph-Aware Deep Point Cloud Compression

### v1

@@ I Introduction
In recent years, robotic perception has shifted from image‐based representations to rich, three‐dimensional point clouds, driven by the need for greater robustness in complex, unstructured and ill-illuminated environments. At the same time, multi‐robot systems are proliferating, and both centralized coordination and peer‐to‐peer collaboration increasingly rely on transmitting 3D data to edge or cloud servers for joint processing [1]. The advent of high‐bandwidth 5G networks and edge‐computing platforms promises lower latency and higher throughput [2], yet the sheer volume and irregular structure of point clouds still impose prohibitive demands on storage, transmission, and real‐time responsiveness.
Point clouds are fundamentally large since each sweep can contain thousands of spatial samples and they lack the regular grid structure that makes images compliant to classical compression schemes. Under realistic bandwidth constraints and with intermittent connections, naive streaming of full‐resolution clouds leads to dropped frames, degraded map accuracy [3], and ultimately failures in perception‐driven tasks such as obstacle avoidance or cooperative mapping [4]. While semantic scene graphs have proven effective as compact, task‐oriented abstractions for navigation and planning existing compression pipelines rarely exploit their structural information to guide lossy encoding of raw geometry.
@@
While previous works demonstrate the advantages of embedding relational and semantic information in 3D representations, they focus on mapping, localization [5], or scene synthesis [6, 7] rather than data compression. To our knowledge, no existing method has systematically fused graph- or structure-aware abstractions with point cloud compression. In this letter, we close this gap by introducing a scene graph–conditioned autoencoder for point cloud compression. Our framework learns compact latent vectors that jointly preserve geometric fidelity and relational structure, yielding a unified representation that not only enables highly efficient compression but could also further support downstream tasks such as semantic scene understanding, place recognition and beyond.
@@
In sum, the main contributions of this letter are: (a) A graph‐conditioned autoencoder for deep point cloud compression. We introduce a novel framework that decomposes raw LiDAR scans into semantically coherent patches using a semantic scene graph, and encodes each patch via a FiLM‐conditioned transformer to jointly compress geometric and semantic cues. (b) Integration of relational structure into compression. Our method leverages semantic scene graphs throughout the pipeline, using semantic conditioning in the encoder and attribute-guided folding in the decoder to preserve relational consistency in the compressed representation. (c) Extensive experiments on SemanticKITTI and nuScenes demonstrate state-of-the-art performance, achieving up to 98% data reduction while preserving both geometric and semantic fidelity. Our method outperforms leading codecs such as Draco [8] and the MPEG approach by Mekuria et al. [9], and maintains strong performance in downstream robotic tasks such as multi-agent pose graph optimization and map merging, even under strict bandwidth constraints.
@@ II Related Work
Point Cloud Compression: Classical methods based on spatial partitioning, such as Octrees [10], provide compact representations by recursively subdividing space into occupied and free regions. While effective at reducing memory usage, they often fail to preserve fine-grained geometry at high compression ratios. Beyond this, several approaches exploit structural redundancies. Sun et al. [11] proposed a clustering-based method that segments range images into ground and object regions, followed by region-specific predictive encoding. Feng et al. [12] introduced a spatio-temporal compression framework combining keyframe selection and iterative plane fitting to model static structures across sequences.
@@
More recently, learning-based approaches have further enhanced compression performance. Wiesmann et al. [13] presented a deep convolutional autoencoder that directly operates on dense point cloud maps, using Kernel Point Convolutions (KPConv [14]) for feature extraction and a deconvolution operator for flexible upsampling. Voxel-based and sparse tensor methods have also gained traction, with examples like VoxelContext-Net [15], that augment octree encoding with voxel context modeling through deep entropy estimation, or SparsePCGC [16] which introduced a multiscale sparse tensor framework focused on the most probable occupied voxels.
@@
Regarding range image-based methods, FLiCR [17] compresses 3D LiDAR data by projecting it into 2D range images and applying lossy quantization and downsampling, enabling lightweight real-time operation on resource-constrained platforms. RIDDLE [18] improves compression by modeling local differences between range image pixels through deep delta encoding, significantly reducing redundancy while preserving fine structural details. Moreover, You et al. proposed RENO [19], a real-time neural compression framework that leverages sparse tensor representations for efficient geometry encoding, achieving high compression efficiency while maintaining fast inference suitable for time-critical applications. Finally, semantic-aware compression methods [20] have also emerged, with scene priors integrated into the compression pipeline, jointly optimizing for geometric fidelity and downstream perception tasks.
@@
While most compression approaches focus on geometric preservation and memory reduction, some recent works explore compression jointly with downstream applications. RecNet [21] proposes an invertible point cloud encoding via range image embeddings, enabling efficient multi-robot map sharing, reconstruction and place recognition. This trend towards combining compression with task-driven objectives highlights the potential of designing representations that are not only compact but also structurally and semantically informative.
@@
Graph– and Structure–Aware Representations: Beyond compression, several works focus on enriching point cloud representations with structural, relational, and semantic information. SegMap [5] introduced a segment-based mapping framework, partitioning 3D LiDAR point clouds into local segments and encoding them with compact learned descriptors. While it demonstrates the benefits of structure-aware feature learning for tasks such as global localization and 3D reconstruction, it primarily focuses on segment-level representations and does not explicitly model the relational structure between objects. Other works pursued explicit scene graph construction from 3D data. Armeni et al. [6] proposed 3D Scene Graphs that organize environments into hierarchical semantic and geometric structures, while Gao et al. [7] developed hierarchical scene graphs for indoor scene synthesis, highlighting the role of relational context in modeling complex spaces.
@@
At the network architecture level, point cloud processing has evolved toward dynamic, graph-based methods. DGCNN [22] introduced dynamic edge convolution operations that adaptively capture local geometric structures during learning, while RandLA-Net [23] proposed lightweight local feature aggregation schemes for large-scale outdoor scenes. These advances emphasize the advantages of relational inductive biases for 3D understanding, while object-centric mapping approaches further demonstrate the value of structure-aware modeling.
@@ III The Proposed Framework
Our goal is to compress a point cloud into compact latent representations, significantly reducing data size, while preserving structural and semantic integrity for efficient transmission between robotic agents. Given a point cloud $\mathcal{P}\subset\mathbb{R}^{3}$, we first apply a semantic segmentation mapping $S:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$, yielding a labeled point cloud $\bar{\mathcal{P}}=S(\mathcal{P})\subset\mathbb{R}^{4}$. We then construct a semantic scene graph, decomposing the cloud into semantically coherent clusters (nodes). Each node is then independently encoded by a semantic-aware encoder conditioned via FiLM [24], generating compact latent vectors. These latent vectors, combined with the scene graph, guide a folding-based decoder to accurately reconstruct the original point cloud.
@@ III-B2 Encoder
Semantic Conditioning via FiLM: To inject semantic context, we embed the node-level semantic label $l_{i}$ using an Embedding layer $E_{\theta}^{s}$, producing a semantic descriptor: $\boldsymbol{s}_{i}=E_{\theta}^{s}(l_{i}),\>\>\text{where}\>\>\boldsymbol{s}_{i}\in\mathbb{R}^{D_{s}}.$ We condition the per-point features using Feature-wise Linear Modulation (FiLM) [24], which applies a learned affine transformation to each feature dimension: $\boldsymbol{h}_{j}^{\prime}=\gamma(\boldsymbol{s}_{i})\odot\boldsymbol{h}_{j}+\beta(\boldsymbol{s}_{i}),$ where $\gamma(\cdot)$ and $\beta(\cdot)$ are MLPs that output scaling and shifting vectors respectively, and $\odot$ denotes element-wise multiplication. This conditioning enables the encoder to adapt its representation based on semantic class, improving generalization across heterogeneous patches.
@@
The Multi-Head Self-Attention module [25] computes attention across all point features, enabling rich contextual encoding of intra-patch geometric relationships. Residual connections and dropout are applied at both attention and Feed-Forward stages to improve stability and generalization.
@@ III-B3 Decoder
Coarse Point Generation: We begin by generating $M$ coarse patch points $\hat{\pi}_{i}^{\text{init}}=\{\boldsymbol{x}_{m}^{\text{init}}\}$ within the node’s bounding box, where each point is sampled using a uniform distribution $\mathcal{U}\big(\boldsymbol{c}_{i},\boldsymbol{e}_{i},\boldsymbol{R}_{i}\big).$ Next, the latent vector $\boldsymbol{z}_{i}$ is mapped to coarse point features via a feed-forward network comprising a two-layer MLP with ReLU activations: $\boldsymbol{f}_{i}^{{c}}=\Phi_{\theta}^{{c}}(\boldsymbol{z}_{i}),\,\boldsymbol{f}_{i}^{{c}}\in\mathbb{R}^{M\times D_{f}^{c}},$ where $D_{f}^{c}$ denotes the dimensionality of the associated coarse features, and $\boldsymbol{f}_{i}^{c}=\{\boldsymbol{f}_{m}^{c}\}_{m=1}^{M}$ with each $\boldsymbol{f}_{m}^{c}\in\mathbb{R}^{D_{f}^{c}}$ representing the feature of the $m$-th coarse point. These features are then processed by a separate MLP: $\Delta\boldsymbol{x}_{m}^{c}=\texttt{MLP}^{\text{offset}}_{\theta}(\mathbf{f}_{m}^{c}),\,m=1,\dots,M,\,\Delta\boldsymbol{x}_{m}^{c}\in\mathbb{R}^{3},$ allowing the model to learn the rough shape of each patch. This uniform initialization provides a structured set of base coordinates, while the latent-conditioned offsets adapt these points to the true geometry, ensuring stable training and faithful reconstruction. The coarse patch points $\hat{\pi}_{i}^{c}=\{\boldsymbol{x}_{m}^{c}\}\in\mathbb{R}^{M\times 3}$ are formed as: $\hat{\pi}_{i}^{c}=\big\{\boldsymbol{x}_{m}^{\text{init}}+\Delta\boldsymbol{x}_{m}^{c}\mid\boldsymbol{x}_{m}^{\text{init}}\in\hat{\pi}_{i}^{\text{init}};\;m=1,\dots,M\big\}.$ Additionally, we predict a coarse confidence mask $\hat{\mu}_{i}^{c}$ indicating the validity of each coarse point via a small MLP followed by a sigmoid activation function $\sigma$: $\hat{\mu}_{i}^{c}=\sigma\left(\texttt{MLP}_{\theta}^{\text{mask}}(\boldsymbol{f}_{i}^{c})\right),\,\hat{\mu}_{i}^{c}\in[0,1]^{M}.$ This coarse patch defines the initial structure upon which the fine reconstruction is built.
@@
Fine Reconstruction via Folding: To produce a higher resolution output of size $N$, we employ a folding-based upsampling strategy [26] that refines each of the $M$ coarse points using a structured 2D grid of $G$ positions, such that $N=M\cdot G$. Each coarse point is thus expanded into a local neighborhood of fine points through learned geometric offsets. Given the set of coarse patch points $\hat{\pi}_{i}^{c}$ and their associated features $\boldsymbol{f}_{i}^{c}=\{\boldsymbol{f}_{m}^{c}\}$, we define a folding grid $\boldsymbol{G}(\hat{\pi}_{i}^{c})=\{\boldsymbol{g}_{n}\}_{n=1}^{g\times g}\in\mathbb{R}^{2}$, where $G=g^{2}$ and each $\boldsymbol{g}_{n}$ is a 2D coordinate. For each pair of coarse features and grid locations, a shared feed-forward network predicts a 3D offset: $\Delta\boldsymbol{x}_{m,n}=\Phi_{\theta}^{\text{fold}}(\boldsymbol{f}_{m}^{c},\boldsymbol{g}_{n}),\,\Delta\boldsymbol{x}_{m,n}\in\mathbb{R}^{3},$ where $m=1,\dots,M$ and $n=1,\dots,G$. Then, we tile the grid $\boldsymbol{G}$ around each coarse point by defining base coordinates: $\tilde{\boldsymbol{x}}_{m,n}=\boldsymbol{x}_{m}^{c},\,\forall\,\,n=1,\dots,G,$ and obtain the final fine-level reconstruction by adding the learned offsets: $\hat{\pi}_{i}=\big\{\tilde{\boldsymbol{x}}_{m,n}+\Delta\boldsymbol{x}_{m,n}\mid m=1,\dots,M;\;n=1,\dots,G\big\}$. Finally, we predict the per-point confidence mask $\hat{\mu}_{i}$ using another MLP network, conditioned on the coarse features and their associated confidence scores: $\hat{\mu}_{i}=\sigma(\texttt{MLP}_{\phi}^{\text{mask}}(\boldsymbol{f}_{i}^{c},\hat{\mu}_{i}^{c})),\,\hat{\mu}_{i}\in[0,1]^{N}$. This upsampling process offers a resolution-aware reconstruction of fine-grained geometries, while maintaining semantic coherence through the structured latent representation and bounding box context.

### v2

@@ I Introduction
In recent years, robotic perception has shifted from image‐based representations to rich, three‐dimensional point clouds, driven by the need for greater robustness in complex, unstructured and ill-illuminated environments. At the same time, multi‐robot systems are proliferating, and both centralized coordination and peer‐to‐peer collaboration increasingly rely on transmitting 3D data to edge or cloud servers for joint processing [seisa2022edge]. The advent of high‐bandwidth 5G networks and edge‐computing platforms promises lower latency and higher throughput [damigos2024-5G], yet the sheer volume and irregular structure of point clouds still impose prohibitive demands on storage, transmission, and real‐time responsiveness.
Point clouds are fundamentally large since each sweep can contain thousands of spatial samples and they lack the regular grid structure that makes images compliant to classical compression schemes. Under realistic bandwidth constraints and with intermittent connections, naive streaming of full‐resolution clouds leads to dropped frames, degraded map accuracy [damigos2023com_aware], and ultimately failures in perception‐driven tasks such as obstacle avoidance or cooperative mapping [stathoulopoulos2024frame]. While semantic scene graphs have proven effective as compact, task‐oriented abstractions for navigation and planning existing compression pipelines rarely exploit their structural information to guide lossy encoding of raw geometry.
@@
While previous works demonstrate the advantages of embedding relational and semantic information in 3D representations, they focus on mapping, localization [dube2020segmap], or scene synthesis [armeni20193d, gao2023scenehgn] rather than data compression. To our knowledge, no existing method has systematically fused graph- or structure-aware abstractions with point cloud compression. In this letter, we close this gap by introducing a scene graph–conditioned autoencoder for point cloud compression. Our framework learns compact latent vectors that jointly preserve geometric fidelity and relational structure, yielding a unified representation that not only enables highly efficient compression but could also further support downstream tasks such as semantic scene understanding, place recognition and beyond.
@@
In sum, the main contributions of this letter are: (a) A graph‐conditioned autoencoder for deep point cloud compression. We introduce a novel framework that decomposes raw LiDAR scans into semantically coherent patches using a semantic scene graph, and encodes each patch via a FiLM‐conditioned transformer to jointly compress geometric and semantic cues. (b) Integration of relational structure into compression. Our method leverages semantic scene graphs throughout the pipeline, using semantic conditioning in the encoder and attribute-guided folding in the decoder to preserve relational consistency in the compressed representation. (c) Extensive experiments on SemanticKITTI and nuScenes demonstrate state-of-the-art performance, achieving up to 98% data reduction while preserving both geometric and semantic fidelity. Our method outperforms leading codecs such as Draco [draco] and the MPEG approach by Mekuria et al. [mekuria2017mpeg], and maintains strong performance in downstream robotic tasks such as multi-agent pose graph optimization and map merging, even under strict bandwidth constraints.
@@ II Related Work
Point Cloud Compression: Classical methods based on spatial partitioning, such as Octrees [octattention], provide compact representations by recursively subdividing space into occupied and free regions. While effective at reducing memory usage, they often fail to preserve fine-grained geometry at high compression ratios. Beyond this, several approaches exploit structural redundancies. Sun et al. [Sun2019clustering] proposed a clustering-based method that segments range images into ground and object regions, followed by region-specific predictive encoding. Feng et al. [Feng2020] introduced a spatio-temporal compression framework combining keyframe selection and iterative plane fitting to model static structures across sequences.
@@
More recently, learning-based approaches have further enhanced compression performance. Wiesmann et al. [wiesmann2021deep] presented a deep convolutional autoencoder that directly operates on dense point cloud maps, using Kernel Point Convolutions (KPConv [thomas2019KPConv]) for feature extraction and a deconvolution operator for flexible upsampling. Voxel-based and sparse tensor methods have also gained traction, with examples like VoxelContext-Net [Que2021VoxelContext], that augment octree encoding with voxel context modeling through deep entropy estimation, or SparsePCGC [Wang2023tensor] which introduced a multiscale sparse tensor framework focused on the most probable occupied voxels.
@@
Regarding range image-based methods, FLiCR [Heo2022FLiCR] compresses 3D LiDAR data by projecting it into 2D range images and applying lossy quantization and downsampling, enabling lightweight real-time operation on resource-constrained platforms. RIDDLE [Theis2022riddle] improves compression by modeling local differences between range image pixels through deep delta encoding, significantly reducing redundancy while preserving fine structural details. Moreover, You et al. proposed RENO [you2025reno], a real-time neural compression framework that leverages sparse tensor representations for efficient geometry encoding, achieving high compression efficiency while maintaining fast inference suitable for time-critical applications. Finally, semantic-aware compression methods [Zhao2022semantic] have also emerged, with scene priors integrated into the compression pipeline, jointly optimizing for geometric fidelity and downstream perception tasks.
@@
While most compression approaches focus on geometric preservation and memory reduction, some recent works explore compression jointly with downstream applications. RecNet [stathoulopoulos2024recnet] proposes an invertible point cloud encoding via range image embeddings, enabling efficient multi-robot map sharing, reconstruction and place recognition. This trend towards combining compression with task-driven objectives highlights the potential of designing representations that are not only compact but also structurally and semantically informative.
@@
Graph– and Structure–Aware Representations: Beyond compression, several works focus on enriching point cloud representations with structural, relational, and semantic information. SegMap [dube2020segmap] introduced a segment-based mapping framework, partitioning 3D LiDAR point clouds into local segments and encoding them with compact learned descriptors. While it demonstrates the benefits of structure-aware feature learning for tasks such as global localization and 3D reconstruction, it primarily focuses on segment-level representations and does not explicitly model the relational structure between objects. Other works pursued explicit scene graph construction from 3D data. Armeni et al. [armeni20193d] proposed 3D Scene Graphs that organize environments into hierarchical semantic and geometric structures, while Gao et al. [gao2023scenehgn] developed hierarchical scene graphs for indoor scene synthesis, highlighting the role of relational context in modeling complex spaces.
@@
At the network architecture level, point cloud processing has evolved toward dynamic, graph-based methods. DGCNN [wang2019dynamic] introduced dynamic edge convolution operations that adaptively capture local geometric structures during learning, while RandLA-Net [hu2020randla] proposed lightweight local feature aggregation schemes for large-scale outdoor scenes. These advances emphasize the advantages of relational inductive biases for 3D understanding, while object-centric mapping approaches further demonstrate the value of structure-aware modeling.
@@ III The Proposed Framework
Our goal is to compress a point cloud into compact latent representations, significantly reducing data size, while preserving structural and semantic integrity for efficient transmission between robotic agents. Given a point cloud $\mathcal{P}\subset\mathbb{R}^{3}$, we first apply a semantic segmentation mapping $S:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$, yielding a labeled point cloud $\bar{\mathcal{P}}=S(\mathcal{P})\subset\mathbb{R}^{4}$. We then construct a semantic scene graph, decomposing the cloud into semantically coherent clusters (nodes). Each node is then independently encoded by a semantic-aware encoder conditioned via FiLM [perez2018film], generating compact latent vectors. These latent vectors, combined with the scene graph, guide a folding-based decoder to accurately reconstruct the original point cloud.
@@ III-B2 Encoder
Semantic Conditioning via FiLM: To inject semantic context, we embed the node-level semantic label $l_{i}$ using an Embedding layer $E_{\theta}^{s}$, producing a semantic descriptor: $\boldsymbol{s}_{i}=E_{\theta}^{s}(l_{i}),\>\>\text{where}\>\>\boldsymbol{s}_{i}\in\mathbb{R}^{D_{s}}.$ We condition the per-point features using Feature-wise Linear Modulation (FiLM) [perez2018film], which applies a learned affine transformation to each feature dimension: $\boldsymbol{h}_{j}^{\prime}=\gamma(\boldsymbol{s}_{i})\odot\boldsymbol{h}_{j}+\beta(\boldsymbol{s}_{i}),$ where $\gamma(\cdot)$ and $\beta(\cdot)$ are MLPs that output scaling and shifting vectors respectively, and $\odot$ denotes element-wise multiplication. This conditioning enables the encoder to adapt its representation based on semantic class, improving generalization across heterogeneous patches.
@@
The Multi-Head Self-Attention module [attentionTrans] computes attention across all point features, enabling rich contextual encoding of intra-patch geometric relationships. Residual connections and dropout are applied at both attention and Feed-Forward stages to improve stability and generalization.
@@ III-B3 Decoder
Coarse Point Generation: We begin by generating $M$ coarse patch points $\hat{\pi}_{i}^{\text{init}}=\{\boldsymbol{x}_{m}^{\text{init}}\}$ within the node’s bounding box, where each point is sampled using a uniform distribution $\mathcal{U}\big(\boldsymbol{c}_{i},\boldsymbol{e}_{i},\boldsymbol{R}_{i}\big).$ Next, the latent vector $\boldsymbol{z}_{i}$ is mapped to coarse point features via a feed-forward network comprising a two-layer MLP with ReLU activations: $\boldsymbol{f}_{i}^{{c}}=\Phi_{\theta}^{{c}}(\boldsymbol{z}_{i}),\,\boldsymbol{f}_{i}^{{c}}\in\mathbb{R}^{M\times D_{f}^{c}},$ where $D_{f}^{c}$ denotes the dimensionality of the associated coarse features, and $\boldsymbol{f}_{i}^{c}=\{\boldsymbol{f}_{m}^{c}\}_{m=1}^{M}$ with each $\boldsymbol{f}_{m}^{c}\in\mathbb{R}^{D_{f}^{c}}$ representing the feature of the $m$-th coarse point. These features are then processed by a separate MLP: $\Delta\boldsymbol{x}_{m}^{c}=\texttt{MLP}^{\text{offset}}_{\theta}(\boldsymbol{f}_{m}^{c}),\,m=1,\dots,M,\,\Delta\boldsymbol{x}_{m}^{c}\in\mathbb{R}^{3},$ allowing the model to learn the rough shape of each patch. This uniform initialization provides a structured set of base coordinates, while the latent-conditioned offsets adapt these points to the true geometry, ensuring stable training and faithful reconstruction. The coarse patch points $\hat{\pi}_{i}^{c}=\{\boldsymbol{x}_{m}^{c}\}\in\mathbb{R}^{M\times 3}$ are formed as: $\hat{\pi}_{i}^{c}=\big\{\boldsymbol{x}_{m}^{\text{init}}+\Delta\boldsymbol{x}_{m}^{c}\mid\boldsymbol{x}_{m}^{\text{init}}\in\hat{\pi}_{i}^{\text{init}};\;m=1,\dots,M\big\}.$ Additionally, we predict a coarse confidence mask $\hat{\mu}_{i}^{c}$ indicating the validity of each coarse point via a small MLP followed by a sigmoid activation function $\sigma$: $\hat{\mu}_{i}^{c}=\sigma\left(\texttt{MLP}_{\theta}^{\text{mask}}(\boldsymbol{f}_{i}^{c})\right),\,\hat{\mu}_{i}^{c}\in[0,1]^{M}.$ This coarse patch defines the initial structure upon which the fine reconstruction is built.
@@
Fine Reconstruction via Folding: To produce a higher resolution output of size $N$, we employ a folding-based upsampling strategy [foldingnet] that refines each of the $M$ coarse points using a structured 2D grid of $G$ positions, such that $N=M\cdot G$. Each coarse point is thus expanded into a local neighborhood of fine points through learned geometric offsets. Given the set of coarse patch points $\hat{\pi}_{i}^{c}$ and their associated features $\boldsymbol{f}_{i}^{c}=\{\boldsymbol{f}_{m}^{c}\}$, we define a folding grid $\boldsymbol{G}(\hat{\pi}_{i}^{c})=\{\boldsymbol{g}_{n}\}_{n=1}^{g\times g}\in\mathbb{R}^{2}$, where $G=g^{2}$ and each $\boldsymbol{g}_{n}$ is a 2D coordinate. For each pair of coarse features and grid locations, a shared feed-forward network predicts a 3D offset: $\Delta\boldsymbol{x}_{m,n}=\Phi_{\theta}^{\text{fold}}(\boldsymbol{f}_{m}^{c},\boldsymbol{g}_{n}),\,\Delta\boldsymbol{x}_{m,n}\in\mathbb{R}^{3},$ where $m=1,\dots,M$ and $n=1,\dots,G$. Then, we tile the grid $\boldsymbol{G}$ around each coarse point by defining base coordinates: $\tilde{\boldsymbol{x}}_{m,n}=\boldsymbol{x}_{m}^{c},\,\forall\,\,n=1,\dots,G,$ and obtain the final fine-level reconstruction by adding the learned offsets: $\hat{\pi}_{i}=\big\{\tilde{\boldsymbol{x}}_{m,n}+\Delta\boldsymbol{x}_{m,n}\mid m=1,\dots,M;\;n=1,\dots,G\big\}$. Finally, we predict the per-point confidence mask $\hat{\mu}_{i}$ using another MLP network, conditioned on the coarse features and their associated confidence scores: $\hat{\mu}_{i}=\sigma(\texttt{MLP}_{\phi}^{\text{mask}}(\boldsymbol{f}_{i}^{c},\hat{\mu}_{i}^{c})),\,\hat{\mu}_{i}\in[0,1]^{N}$. This upsampling process offers a resolution-aware reconstruction of fine-grained geometries, while maintaining semantic coherence through the structured latent representation and bounding box context.